# Can dissipation induce a transition between many-body localized and thermal states?

### v1

@@ III Discussion on the effects of system size
When the distribution of single-particle excitation is extended or localized, the corresponding SIPR respectively tends to 00 or a finite nonzero value. The SIPR corresponding to the ground state and the highest excited state are $0.023$ (close to 00) and $0.301$, respectively [Fig. 4], further indicating that they are extended and localized, respectively.
Figure 4: The distribution of $\delta n$ at each lattice site after adding a particle to a system with $L=144$ and $N=32$ for the ground state (orange dots) and the highest excited state (blue dots), with their corresponding SIPR values being $0.023$ and $0.301$, respectively.
@@
IV Conclusion
We have investigated the impact of a type of bond dissipation on the quasiperiodic $t_{1}-t_{2}$ model with NN interactions, which possesses a mobility edge in the absence of interactions. By calculating the level spacing distribution, we observed the potential existence of a MBME. We further analyzed the distribution of the steady-state density matrix and revealed that dissipation can drive the many-body system into specific states primarily located in either the thermal region or the MBL region, regardless of the initial states. Thus, dissipation can be used to induce transitions between thermal and MBL states, allowing for the control of particle transport behaviors, as demonstrated by our analysis of the system’s transport properties after removing dissipation. Finally, we used DMRG to study single-particle excitations in the ground state and highest excited state for large system sizes, demonstrating their different localization behaviors. Our findings suggest that dissipation can trigger transitions between thermal and MBL states, offering a novel method for experimentally identifying the presence of the MBME.
Acknowledgements
@@
Appendix A Change in current: An example using the AA model
We take the interacting Aubry-André (AA) model (i.e., with next-nearest-neighbor hopping strength $t_{2}=0$ in the Hamiltonian of Eq. (1)) as an example to illustrate the different current changes between thermal and many-body localized states after introducing an electric field. We still fix one-third filling. We consider only the ground state and select two extreme values of the quasiperiodic potential: $\lambda=0.1$, representing the thermal state property, and $\lambda=10$, representing the MBL state property.

### v2

@@ II.3 Discussion on the effects of system size
If the ground state or the highest excited state is thermal, this added particle density profile should be extended, and therefore the SIPR should approach 00. However, if the ground state or the highest excited state is many-body localized, the added particle’s effect on the system does not propagate to distant regions, meaning this effect is localized. In this case, the added particle density profile should be localized, and the corresponding SIPR will not approach 00 but instead settle at a finite value. We add a particle to a system with size $L=144$ and particle number $N=32$, and show the distribution of $\delta n$ at each lattice site in Fig. 4(a). We see that for the ground state (the orange dots), the single particle excitation is distributed quite uniformly across different lattice sites, indicating an extended property. In contrast, for the highest excited state (the blue dots), $\delta n_{j}$ is primarily concentrated on a few lattice sites, showing localized behavior. Fig. 4(b) shows the variation of the SIPR with system size. We observe that for the ground state, the SIPR approaches 00 as the size increases, while for the highest excited state, the SIPR eventually stabilizes at a certain value and does not decrease with increasing size. These results correspond to those in Fig. 1, indicating a significant difference in localization properties between the ground state and the highest excited state in larger systems. We note that the ground state and the highest excited state exhibit different localization properties, which, apart from the possible existence of a MBME, could also arise from boundary effects in finite-size systems, topological effects (such as the presence of topological boundary states or defects), electron-phonon coupling, or irregular band structures, among others. However, in the context of the system we study-where a ME is present in both single-particle and small-size many-body systems-the most likely explanation lies in the existence of a MBME. This interpretation is consistent with recent experiments DeteME1 ; DeteME2 that observed distinct localization behaviors of the ground and highest excited states as evidence for the existence of a ME. While we do not claim to numerically prove the existence of a MBME, this approach may offer useful insights into addressing finite-size effects in its investigation.
In the DMRG calculation, we always retain $40$ states when truncating the reduced density matrix. Changing the number of retained states does not significantly affect the $\delta n_{j}$ and SIPR for the highest excited state, as its entanglement satisfies the area law, and DMRG can provide accurate results. However, for the ground state, changing the number of retained states during the calculation causes noticeable changes in $\delta n_{j}$, which indicates that the DMRG results for the thermal state are not precise. Nevertheless, $\delta n_{j}$ remains delocalized, and the SIPR does not change significantly with the number of retained states (see Appendix D). Therefore, while the DMRG results for the ground state are not accurate, they are sufficient to demonstrate its delocalized properties.
II.4 Discussion on detecting the MBME
@@
III Conclusion
We have investigated the impact of a type of bond dissipation on the quasiperiodic $t_{1}-t_{2}$ model with NN interactions, which possesses a mobility edge in the absence of interactions. By calculating the level spacing distribution, we observed signatures consistent with the presence of a MBME. We further analyzed the distribution of the steady-state density matrix and revealed that dissipation can drive the many-body system into specific states primarily located in either the thermal region or the MBL region, regardless of the initial states. Thus, dissipation can be used to induce transitions between thermal and MBL states, allowing for the control of particle transport behaviors, as demonstrated by our analysis of the system’s transport properties after removing dissipation. We note that bond dissipation only needs to act for a short period of time to drive the system into a steady state, after which it can be removed. Moreover, even substantial fluctuations in the strength of the bond dissipation do not alter the steady-state results. Therefore, the physical effects induced by this dissipation are highly stable. Finally, we used DMRG to study single-particle excitations in the ground state and highest excited state for large system sizes, demonstrating their different localization behaviors, which implies that MBME may exist at large sizes. Therefore, the dissipation-induced thermal-MBL transition should also exist at large sizes, providing a novel method for experimentally identifying the presence of the MBME.
Acknowledgements.
@@
Appendix A Compute the steady state
To numerically compute the steady state, we represent the density matrix $\rho$ as a vector and the Liouvillian superoperator $\mathcal{L}$ as a matrix. In the computational basis $\{\left|{k}\right\rangle\}$, the density matrix can be expressed as $\rho=\sum_{j,k}\rho_{jk}\left|{j}\right\rangle\left\langle{k}\right|$, which is then mapped to a vector representation as:
$\left|{\rho}\right\rangle=\sum_{j,k}\rho_{jk}\left|{j}\right\rangle\otimes\left|{k}\right\rangle.$ (8)
Consequently, the Liouvillian superoperator $\mathcal{L}$ can be represented as a matrix:
$\mathcal{L}$ $=$ $\sum_{j}\left[{O}_{j}\otimes{O}_{j}^{*}-\frac{1}{2}\left({O}_{j}^{\dagger}{O}_{j}\otimes\mathbbm{1}-\mathbbm{1}\otimes{O}_{j}^{\mathrm{T}}{O}_{j}^{*}\right)\right]$ (9)
$-i\left({H}\otimes\mathbbm{1}-\mathbbm{1}\otimes{H}^{\mathrm{T}}\right).$
We diagonalize the matrix representation of $\mathcal{L}$ to obtain its eigenvalues and eigenvectors $\left|{\rho}\right\rangle$. The eigenvalue of 00 corresponds to the steady state $\left|{\rho_{s}}\right\rangle$. Using the mapping defined in (8), we can reconstruct the steady-state density matrix $\rho_{s}$. For a Hamiltonian with a Hilbert space dimension $D_{H}$, the Liouvillian $\mathcal{L}$, as described in Eq. (9), is a matrix of size $D_{H}^{2}\times{}D_{H}^{2}$. Therefore, for the many-body system with dissipation, we only compute up to $L=9$ in the main text.
Appendix B Average IPR for different energy windows
The localization properties of different energy windows shown in Fig. 1(b) of the main text can also be characterized using the inverse participation ratio (IPR), which is defined as ${\rm IPR}=\sum_{j}|\phi_{j}|^{4}$, where $\phi_{j}$ is the coefficient of the eigenstate $|\psi_{n}\rangle$ in the computational basis $\{|j\rangle\}$, expressed as $\phi_{j}=\langle j|\psi_{n}\rangle$. As the size of the Hilbert space increases, the ${\rm IPR}$ for thermal states tends to zero, while for MBL states, the ${\rm IPR}$ remains a finite nonzero value. Fig. 5 shows the average ${\rm IPR}$ within each energy window. Similar to the results of the level statistics in Fig. 1 of the main text, as the quasiperiodic potential strength $\lambda$ increases, the states in the highest energy window localize first.
Figure 5: Variation of $\left\langle{\rm IPR}\right\rangle$ as a function of energy density ($\epsilon$) and quasiperiodic potential strength $\lambda$, using the same parameters as in Fig. 1(b) of the main text.
Appendix C Current Changes After Introducing an Electric Field in the Interacting AA Model
In this section, we take the interacting Aubry-André (AA) model ($H=-\sum_{j}\left(t_{1}c_{j+1}^{{\dagger}}c_{j}+\text{H.c.}\right)+U\sum_{j}n_{j}n_{j+1}+2\lambda\sum_{j}\cos\left(2\pi\omega j+\phi\right)n_{j}$, i.e., with next-nearest-neighbor hopping strength $t_{2}=0$ in the Hamiltonian of Eq. (1) of the main text) as an example to illustrate the different current changes between thermal and many-body localized states after introducing an electric field. We still fix one-third filling. We consider only the ground state and select two extreme values of the quasiperiodic potential: $\lambda=0.1$, representing the thermal state property, and $\lambda=10$, representing the MBL state property.
@@ Appendix C Current Changes After Introducing an Electric Field in the Interacting AA Model
Appendix D The Impact of the Number of Retained States on SIPR in DMRG Calculations
For thermal states, their entanglement does not follow the area law. Therefore, it is necessary to clarify that the DMRG calculations used in the main text are valid. As mentioned in the main text, although the $\delta n_{j}$ values given by DMRG may change, their delocalized nature remains unchanged. In Fig. 7, we fix $L=48$ and a one-third filling, showing how the SIPR of the ground state and the highest excited state varies with the number of retained states $n_{s}$ in truncating the reduced density matrix. We observe that the SIPR corresponding to the added particle density profile of the thermal state (red data points) does not change with $n_{s}$. This indicates that the SIPR for the thermal state obtained from our DMRG calculations is valid, and using SIPR to reflect localized and delocalized properties is feasible.
Figure 7: SIPR as a function of $n_{s}$ with $L=48$ for the ground state (red dots) and the highest excited state (blue dots). We sweep $20$ cycles in finite size DMRG.